IDENTIFICATION AND ASSIGNMENT OF THE FIRST EXCITED TORSIONAL STATE OF CH_{2}DOH WITHIN THE o_{2}, e_{2}, AND o_{3} TORSIONAL LEVELS

\underline{JOHN C. PEARSON}, SHANSHAN YU, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA; L. H. COUDERT, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France; L. MARGULÈS, R. A. MOTIYENKO, Laboratoire PhLAM, UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; and S. KLEE, Physikalisch-Chemisches Institut, Justus-Liebig-Universität Gießen, 35392 Gießen, Germany.

Theoretical models describing the details of asymmetric-top asymmetric-frame internal rotation remain to be fully developed and tested especially in excited torsional states. The spectrum of CH_{2}DOH offers a unique opportunity to test and develop asymmetric-top asymmetric-frame internal rotation theory. The theoretical energy levels predicted by El Hilali et al.^{a} and combination differences of the 76 assigned torsional subbands coupled with the experimental microwave energy levels of the ground state^{b} served as a basis for assigning the excited torsional state. The existing microwave spectra was supplemented with recordings of 1308–2010 GHz and 2450–2700 GHz. This facilitated extension of the ground state assignments to K = 14 and identified a number of torsional interactions. In this paper we report assignment of the o_{2}, e_{2}, and o_{3} torsional levels to K = 9. All the torsional levels in the first excited torsional state have been connected with microwave accuracy transitions except o_{3} K = 2. Strong rotational interactions between the o_{2} K = 0 and K = 2 states and the e_{2} K = 4, o_{2} K = 1, o_{3} K = 2 and o_{3} K = 3 levels are observed. A week avoided crossing between e_{1} K = 12 and e_{2} K = 8 at J = 20 has also been identified. When the microwave results are combined with the existing infrared assignments it is now possible to predict the entire v_{t} = 0 to v_{t} = 1 torsional band to better than infrared experimental accuracy, greatly simplifying subsequent infrared assignments.

aEl Hilali, Coudert, Konov, and Klee, J. Chem. Phys. 135 (2011) 194309.

bPearson, Yu, and Drouin, J. Mol. Spectrosc. 280 (2012) 119.